Q- one mole of N2O4 at 300 k is kept in a closed container under one atmosphere. It is heated to 600 K when 20% by mass of N2O4(g)  decomposes to NO2(g). The resultant pressure is....????


Plz explain me in detail.....

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
the given equation is
N2O4 ------2NO2
mass of 1 mole of N2O4=92
left mass after conversion=92-.2*92=73.6gm
mole of N2O4=.8mole
mole of NO2=18.4/46=.4 mole

as volume,tempaerature is constant therefore the mole is proportional to pressure

total pressure=.8 +.4=1.2atm